If two block with masses 5kg and4kg connected by a string are placed on a horizontal table and are pulled by a force of 30N calculate the tension in the string if frictional force is negligible.Repeat the problem if the coefficient of kinetic friction between the blocks an the table is 0.3N


1
Expert's answer
2021-04-26T17:10:08-0400

1)


"(5+4)a=30\\T=4a=4\frac{10}{3}=13.3\ N"

2)


"(5+4)(a+0.3(9.8))=30\\T=4a=4(0.393)=1.57\ N"